OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−1, 3) and (1, 7)? y − 7 = 4(x − 1) y − 7 = 2(x − 1) y − 3 = 2(x − 1) y − 3 = 4(x + 1)

OpenStudy (anonymous):

@12irun help

OpenStudy (anonymous):

@sidona005 can you help

OpenStudy (anonymous):

@dominique5

OpenStudy (nirmalnema):

(−1, 3) and (1, 7) 7-3/1+1 =4/2 =2 y-7=2(x-1) y-7=2x-2 2x-y+5=0

OpenStudy (anonymous):

whats the answer ? or is it in there

OpenStudy (anonymous):

@nirmalnema

OpenStudy (nirmalnema):

y − 7 = 2(x − 1)

OpenStudy (anonymous):

its blocking it out with question marks @nirmalnema